OpenStudy (anonymous):

Solving the exponential equation algebraically, e^(2x)-6e^x=-5 I know I have to take a natural log of something, but nothing else.

OpenStudy (anonymous):

I typed it into Wolfram and it said, it was log5 as the answer, but I can't understand how it got to that.

OpenStudy (anonymous):

take e^x = A  hence e^x = 1 &&&& e^x=5 now apply log to get your answer...